# Universality of Rényi Entropy in Conformal Field Theory

### v1

@@ I Introduction and Summary
Recently, a powerful method has been developed to discover more of these universal formulas using the thermal effective theory [4, 5, 6, 7, 8, 9, 10, 11, 12]. Compactifying $d$-dimensional CFT on the thermal circle $\mathbb{S}^{1}_{\beta}$ at temperature $\beta^{-1}$, the CFT in the high temperature limit $\beta^{d-1}\ll{\rm vol}(\Sigma_{d-1})$ is described by a quantum field theory on $\Sigma_{d-1}$ with a mass gap $m_{\text{gap}}=1/\beta$. It is convenient to use the Kaluza-Klein (KK) ansatz for the $d$-dimensional metric $G_{\mu\nu}$ on $\mathbb{S}^{1}_{\beta}\times\Sigma_{d-1}$ as,
@@
In two dimensions, we can use the hot spot idea to derive a result stronger than the combination of eqs. (8) and (9), which is applicable to arbitrary positive $n$ as,
@@
One might expect a similar formula for arbitrary positive $n$ in higher dimensions. However, it turns out that the hot spot idea is not directly applicable in such a case because of the behavior of the effective temperature $\beta^{-1}$ near the boundary of the entanglement domain, as we will explain in Section V. This does not necessarily mean that eq. (12) does not generalize to higher dimensions, but it suggests that the hot spot idea has a certain limitation. We discuss the necessary conditions for the applicability of the hot spot idea and show that it can be used consistently for the computation in [5] and that the universal formula for OPE coefficients derived there using the hot spot idea remains valid.
@@ III Contribution from Boundary of Entanglement Domain
Thus, we obtain in the $\epsilon\to 0$ limit for arbitrary positive $n$,
@@ IV Entanglement Spectrum
This is consistent with the naive formula (LABEL:eq:naive) for the asymptotic spectrum when integrated against a window function supported on $[0,\Lambda]$ in the $\Lambda\to\infty$ limit. A similar result has been obtained in the context of the Cardy formula in Appendix C of [28].
@@ V Applicability of Hot Spot Idea
In two dimensions, the universal formula (12) for the Rényi entropy originates from the $\epsilon\to 0$ limit, rather than the $n\to 0$ limit. That is why the formula is applicable to arbitrary positive $n$ not just in the $n\rightarrow 0$ limit. Can we find a similar formula in higher dimensions?
@@
The fact that the leading contribution to the Rényi entropy for arbitrary positive $n$ can be universally expressed in terms of the central charge, seems to be a unique property in two dimensions. Our result suggests that, in higher dimensions, the Rényi entropy for arbitrary positive $n$ depends strongly on the details of the theory.
@@ VII Discussion
(i) In this paper, we discussed how the thermal effective action is modified in the presence of spatial boundaries and used the result to estimate the Rényi entropy, the entanglement spectrum, and the asymptotic density of boundary states. We expect many other applications of the thermal effective theory to BCFT. An interesting direction may be to use it to generalize the known universality in two dimensions [29, 30] to higher dimensions
(ii) In this paper, we restricted our attention to the case when the boundary of the entanglement domain $A$ is spherical. It would be interesting to see if the universal formulas such as eq. (9) generalize to arbitrary shapes of $A$ as far as it is connected. A possible approach may be to deform the Weyl invariant metric $\hat{g}_{ij}$ and use the thermal effective action to understand the responses of the Rényi entropy and other physical quantities.
(iii) If CFT has a global symmetry $G$, we can turn on the gauge field and use the method developed in [31, 32, 6] to obtain the symmetry resolution of the Rényi entropy [33, 34, 35, 36], as
@@ Acknowledgments
The authors thank David Simmons-Duffin for helpful discussions and comments on the draft. This research was supported in part by the U.S. Department of Energy, Office of Science, Office of High Energy Physics, under Award Number DE-SC0011632 and by the the Walter Burke Institute for Theoretical Physics at Caltech. HO is also supported in part by the Simons Investigator Award (MP-SIP-00005259) and by JSPS Grants-in-Aid for Scientific Research 23K03379. His work was performed in part at the Kavli Institute for the Physics and Mathematics of the Universe at the University of Tokyo, which is supported by the World Premier International Research Center Initiative, MEXT, Japan, at the Kavli Institute for Theoretical Physics (KITP) at the University of California, Santa Barbara, which is supported by grant NSF PHY-2309135, and at the Aspen Center for Physics, which is supported by NSF grant PHY-1607611. YK is also supported by the INAMORI Frontier Program at Kyushu University and JSPS KAKENHI Grant Number 23K20046.

### v2

@@ I Introduction and Summary
[…]
@@
In two dimensions, we can use the hot spot idea to derive a result stronger than the combination of eqs. (8) and (9), which is applicable to arbitrary $n$ as,
@@
One might expect a similar formula for arbitrary $n$ in higher dimensions. However, it turns out that the hot spot idea is not directly applicable in such a case because of the behavior of the effective temperature $\beta^{-1}$ near the boundary of the entanglement domain, as we will explain in Section V. This does not necessarily mean that eq. (12) does not generalize to higher dimensions, but it suggests that the hot spot idea has a certain limitation. We discuss the necessary conditions for the applicability of the hot spot idea and show that it can be used consistently for the computation in [7] and that the universal formula for OPE coefficients derived there using the hot spot idea remains valid.
@@ III Contribution from Boundary of Entanglement Domain
Thus, we obtain in the $\epsilon\to 0$ limit for arbitrary $n$,
@@ IV Entanglement Spectrum
This is consistent with the naive formula (LABEL:eq:naive) for the asymptotic spectrum when integrated against a window function supported on $[0,\Lambda]$ in the $\Lambda\to\infty$ limit. A similar result has been obtained in the context of the Cardy formula in Appendix C of [32] and in [2].
@@ V Applicability of Hot Spot Idea
In two dimensions, the universal formula (12) for the Rényi entropy originates from the $\epsilon\to 0$ limit, rather than the $n\to 0$ limit. That is why the formula is applicable to arbitrary $n$ not just in the $n\rightarrow 0$ limit. Can we find a similar formula in higher dimensions?
@@
The fact that the leading contribution to the Rényi entropy for arbitrary $n$ can be universally expressed in terms of the central charge, seems to be a unique property in two dimensions. Our result suggests that, in higher dimensions, the Rényi entropy for general $n$ depends on the details of the theory (See [33] for related comments in the context of holography).
@@ VII Discussion
  • In this paper, we discussed how the thermal effective action is modified in the presence of spatial boundaries and used the result to estimate the Rényi entropy, the entanglement spectrum, and the asymptotic density of boundary states. We expect many other applications of the thermal effective theory to BCFT. An interesting direction may be to use it to generalize the known universality in two dimensions [34, 35] to higher dimensions.
  • In this paper, we restricted our attention to the case when the boundary of the entanglement domain $A$ is spherical. It would be interesting to see if the universal formulas such as eq. (9) generalize to arbitrary shapes of $A$ as far as it is connected. A possible approach may be to deform the Weyl invariant metric $\hat{g}_{ij}$ and use the thermal effective action to understand the responses of the Rényi entropy and other physical quantities.
  • If CFT has a global symmetry $G$, we can turn on the gauge field and use the method developed in [36, 37, 8] to obtain the symmetry resolution of the Rényi entropy [38, 39, 40, 41], as
@@ Acknowledgments
The authors thank David Simmons-Duffin for helpful discussions and comments on the draft. This research was supported in part by the U.S. Department of Energy, Office of Science, Office of High Energy Physics, under Award Number DE-SC0011632 and by the the Walter Burke Institute for Theoretical Physics at Caltech. HO is also supported in part by the Simons Investigator Award (MP-SIP-00005259) and by JSPS Grants-in-Aid for Scientific Research 23K03379. His work was performed in part at the Kavli Institute for the Physics and Mathematics of the Universe at the University of Tokyo, which is supported by the World Premier International Research Center Initiative, MEXT, Japan, at the Kavli Institute for Theoretical Physics (KITP) at the University of California, Santa Barbara, which is supported by grant NSF PHY-2309135, and at the Aspen Center for Physics, which is supported by NSF grant PHY-1607611. YK is also supported by the INAMORI Frontier Program at Kyushu University and JSPS KAKENHI Grant Number 23K20046. The authors thank Kyushu University Institute for Advanced Study and RIKEN Interdisciplinary Theoretical and Mathematical Sciences Program. Discussions during the “Kyushu IAS-iTHEMS workshop: Non-perturbative methods in QFT” were useful in completing this work.